OpenStudy (anonymous):

Difference between f(-2) and f(x)=0?

OpenStudy (aum):

For example, if f(x) = x - 5 then, f(-2) = -2 - 5 = -7 f(x) = 0 implies x -5 = 0 or x = 5

OpenStudy (anonymous):

Oh btw the equation is f(x)=3x^2-2x+6 I know what f(-2) is, I just plug that in for x. But what about the other one?

OpenStudy (anonymous):

Wait, I didn't get that.

myininaya (myininaya):

f(x)=0 means you set what f(x) is equal to to zero

OpenStudy (aum):

f(x)=3x^2-2x+6 f(x) = 0 implies 3x^2-2x+6 = 0 solve the quadratic for x.

OpenStudy (anonymous):

Ok, so something like this?:  Wait what would I do after this step?